OpenStudy (anonymous):

Cot a - 2 Cot a / Cot a + 2 Cot a = Sin a / 3 Sin a

OpenStudy (freckles):

also I assume this proving identities I think writing in terms of sin and cos might be a good start

OpenStudy (anonymous):

It gets to -cos a/sin a * sin a / 3 cos a ! How to go about then

OpenStudy (freckles):

\[\frac{-\frac{\cos(a)}{\sin(a)}}{3 \frac{\cos(a)}{\sin(a)}} \]

OpenStudy (freckles):

write the division problem as a multiplication prob

OpenStudy (freckles):

\[\frac{-\cos(a)}{\sin(a)} \cdot \frac{\sin(a)}{3 \cos(a)}\]

OpenStudy (freckles):

cancel some things

OpenStudy (freckles):

and also i assume the problem was \[\frac{\cot(a)-2\cot(a)}{\cot(a)+2\cot(a)}\]

OpenStudy (anonymous):

If sin a cancels and also we cancel cos a/ 3cos a we will have only 1/3 ????....yes the problem is the correct one as you typed!!